# Gravitational Atoms from Topological Stars

### v1

@@ Abstract
We study the bound states of a massive scalar field around a topological star, and show that these are strictly normal modes. This yields a genuine gravitational atom, sharply distinguishing horizonless objects from black holes. It is shown that the modes are controlled by the field’s Compton wavelength compared to the size of the star. When the Compton wavelength is large, the field forms a cloud with a hydrogen-like spectrum, while in the opposite regime it localizes along timelike trajectories. When the two scales are comparable the spectrum becomes richer, and we characterize it in detail allowing the field to carry electric charge and Kaluza–Klein momentum.
@@ I Introduction
The existence of event horizons in our universe is a robust prediction of General Relativity (GR), as follows from combining singularity theorems Penrose (1965), cosmic censorship conjectures Penrose (1969) and the genericity of trapped surfaces Schoen and Yau (1983); Christodoulou (2008). Gravitational wave (GW) observatories Abbott et al. (2016, 2021) and long baseline interferometry Akiyama et al. (2019); Abuter et al. (2020) give support to such predictions on the empirical side. However, horizons pose major conceptual challenges such as the formation of spacetime singularities, the entropy problem and the information paradox Bena et al. (2022); Carballo-Rubio et al. (2025), none of them being resolvable by GR on its own. These reasons are strong enough to motivate the exploration of alternatives to the black hole (BH) paradigm, despite the strong theoretical and observational evidence for their existence Cardoso and Pani (2019).
@@
Topological solitons, as we will refer to the aforementioned objects in this work, are dramatically different from BHs near their surface. The latter feature an event horizon, which for an infalling observer is nothing but pure vacuum, yet it marks the point of no return for any such probe - rendering the inner boundary of a BH spacetime perfectly absorbing. Topological solitons, on the other hand, replace the horizon with a smooth end of spacetime - a “cap”, which, furthermore, supports non-negligible horizon-scale microstructure. This results in an inner boundary that is very much the opposite of the BHs of GR - classically the cap is perfectly reflecting and its physics are certainly not of an empty spacetime. Such a dramatic change should surely be expressed in certain properties that topological solitons have or lack in contrast to BHs. Our aim here is to elucidate one such feature.
However, given these stark differences, how is it that topological solitons provide a good fit to the criteria for BH mimickers exhibited earlier? Most of the GW observations up to date are a consequence of the physics at the unstable photon sphere in the exterior of the spacetime, which effectively governs the linear response to perturbations at early times, irrespective of the presence or absence of a horizon Cardoso et al. (2016); Cardoso and Pani (2017). The topological solitons we are interested in all possess such a photon sphere, and, in specific cases, it has indeed been shown that the prompt response as well as the quasinormal modes relative to neutral, massless (or massive but radiative) fields exhibit a spectrum that parallels the one of BHs, although other sectors with new properties also emerge Heidmann et al. (2023a); Bianchi et al. (2023); Dima et al. (2024, 2025b). In addition, particle motion as well as gravitational and electromagnetic fluctuations have been explored in Dima et al. (2024); Bena et al. (2024); Dima et al. (2025b); Melis et al. (2025); Bini and Di Russo (2025a); Bianchi et al. (2025b); Bini and Di Russo (2025b); Di Russo et al. (2025); Bianchi et al. (2025c), giving strong evidence for their linear stability and again demonstrating important similarities with BHs. Finally, with respect to their tidal deformability, TSs are strikingly similar to magnetic BHs, both exhibiting vanishing tidal Love numbers relative to neutral fields, but non-vanishing ones when charged probes are considered, as shown recently in Pereñiguez and Karnickis (2025).
@@
One such aspect that has remained largely unexplored is the bound states of massive fields around compact objects. In the case of BHs, it is well understood that such fields - for example, dark matter candidates - can grow spontaneously, forming clouds that are bound to the hole, a system known as the gravitational atom Damour et al. (1976); Zouros and Eardley (1979); Detweiler (1980a). The mechanism yielding these clouds is the superradiant amplification due to the hole’s rotation, so, by definition, such gravitational atoms are at most quasi-stationary states, and require that the central hole rotates (see Brito et al. (2015a) for a review). These systems have been widely studied in the literature and have been considered to test possible dark matter candidates Arvanitaki et al. (2010); Arvanitaki and Dubovsky (2011), obtain photon mass bounds Pani et al. (2012); Baumann et al. (2019a), and have been proposed as new sources of gravitational Yoshino and Kodama (2014); Brito et al. (2015b); Arvanitaki et al. (2015) and electromagnetic radiation Rosa and Kephart (2018); Ikeda et al. (2019); Chen et al. (2020). Moreover, when studying binaries, the cloud’s presence can affect the orbital dynamics to varying extents, depending on how significant the cloud’s self-gravity is Ferreira et al. (2017); Bar et al. (2019); Amorim et al. (2019); Vicente and Cardoso (2022); Duque et al. (2024); Dyson et al. (2025); Tomaselli et al. (2024a, b, 2023), and tidal effects can enrich further the binary dynamics Baumann et al. (2019b); Berti et al. (2019); Cardoso et al. (2020); Zhang and Yang (2019); Baumann et al. (2020).
@@
In the case of topological solitons, the perfectly reflecting inner boundary condition, resulting from the smooth end of spacetime in their interior, would imply that the states of such massive fields are not only bound at infinity, but also have nowhere to dissipate to at the surface of the soliton. We expect then that topological solitons will display a normal mode spectrum in addition to their QNMs - a property that is unattainable for the BHs of GR. This clear difference in the response to massive perturbations (between BHs and TSs) undoubtedly opens an interesting new avenue for exploration in view of the vast amount of research that has followed from the discovery of superradiant clouds around BHs.
@@
In this work we confirm this expectation in the case of one of the simplest such geometries that is also one of the most widely studied ones - topological stars (TSs) Bah and Heidmann (2021a, c). They are a one-parameter family of static solutions to the Einstein–Maxwell theory in five spacetime dimensions. Far from the star, the geometry is that of a four-dimensional Minkowski space endowed with a compact, circular extra dimension as in the traditional Kaluza–Klein set up. Closer to the star, the geometry is smooth and supported by electromagnetic fluxes. Despite their simplicity, TSs contain the fundamental ingredients of any microstate geometry - a horizonless, smooth end-of-space, supported by electromagnetic flux. We probe them with an electrically charged, massive scalar field minimally coupled to the Einstein–Maxwell theory. We provide a comprehensive characterization of its bound states, that is, the states that are exponentially suppressed at infinity. Among our results, we show that any such mode is strictly stationary, thus proving that TSs are linearly stable under such perturbations with a characteristic normal mode spectrum. In this sense, TSs yield a more genuine notion of a gravitational atom than BHs, since the latter are only quasi-stationary and, in fact, unstable. We show that, in certain regimes, TS gravitational atoms share similarities with their BH counterparts, but are, in general, markedly different. In addition, the KK momentum and electric charge endow the space of states with a rich structure, mostly absent in the BH case, which we analyze in detail. At the end of the paper, we discuss possible formation mechanisms of TS gravitational atoms focusing on potential instabilities of different natures, superradiant versus ergoregion ones.
@@
While our work focuses on the simple TS, we believe that our results for the existence of normal modes should be widely applicable to any horizonless geometry with perfectly reflecting inner boundary conditions. Irrespective of whether deriving a generalized result is possible, we believe that simple topological solitons are excellent, UV-motivated toy models for exploring horizonless physics with the aim of learning valuable lessons for the future. Building a refined as possible, smooth, horizonless geometry that mimics as closely as possible in our observations the famous Kerr BH is currently a hard technical challenge, but no solid theoretical arguments to the contrary are known Chakraborty and Heidmann (2025). In light of this peak into the future, we would like to emphasize another property of topological solitons that is in definite contrast to BHs and requires more attention. As this article began, classically, the end product of gravitational collapse for sufficiently dense energy distributions is always a BH. The dynamical formation of topological solitons is thus necessarily a process that requires physics beyond classical GR.
@@
This paper is structured as follows. First, we summarize our results in Section I.1. Next, in Section II we discuss generic aspects of massive bound states of TSs. This includes proving that bound modes are strictly normal and showing that in the long (short) Compton wavelength regime, modes are hydrogenic (governed by timelike geodesics). In Section III we provide a thorough analysis of TS normal modes, discussing their regime of existence, classifying them, and also computing their spectrum numerically, including the scenarios where the field carries KK momentum and is electrically charged. We conclude in Section IV discussing our results and open questions.
I.1 Summary
@@
Our results fall into two classes: those that are expected to hold in a qualitatively similar way for compact objects other than TS, and those that are TS-specific. We summarize them here. It is convenient to remind the reader that TSs are characterized by two length scales, the TS’s size $r_{b}$ and the length of the compact extra-dimension $R_{y}$.
@@ I.1.1 Generic results
considering neutral probes (no charge or KK momentum), in the limit $\mu r_{b}\gg 1$, we show that solutions are governed by congruences of timelike geodesics. This allows us to identify two broad families of modes, dubbed cloud and inner modes. The former have a large angular momentum $\ell\gg\mu r_{b}$ and are exponentially localized around stable circular geodesics. They cease to exist for angular momenta lower than $\ell\sim\sqrt{3}r_{s}\mu$, where $r_{s}$ is a length fixed by $r_{b}$ and $R_{y}$ and related to the ISCO by $r_{ISCO}=3r_{s}$. The behavior of modes changes abruptly below that threshold and becomes dictated by the inner structure of the geometry – these are the inner modes. The separation between both families, at $\ell\sim\sqrt{3}r_{s}\mu$, is associated with the ISCO and the modes inherit some of its properties. This is illustrated in Figure 1.
@@
in the opposite regime, where $\mu r_{b}\ll 1$, the normal mode spectrum (restricting here to zero-KK modes) becomes hydrogenic,
$\omega_{n}\approx\mu\left(1-\frac{\mu^{2}r_{s}^{2}}{8}\frac{1}{n_{\text{H}}^{2}}\right)\,,\quad\left(n_{\text{H}}=1,2,...\,,\ \ n_{\text{H}}>\ell\right)$ (1)
A similar behavior is also present in the superradiant quasi-bound states of rotating BHs Detweiler (1980b); Baumann et al. (2019a), in the limit $\mu M\ll 1$, where $M$ is the BH’s mass. One difference is the fine structure constant, here $\sim\mu r_{s}$ while $\sim\mu M$ for BHs. However, the most important difference is that in the TS case these modes are truly stationary states, unlike for BHs where they are unstable.
@@ I.1.2 TS-specific results
for probes with no KK momentum (yet possibly charged), the gravitational potential of TSs for bound modes can yield either one or two classically allowed regions where modes can localize. This leads to a precise classification of bound states in TS modes, cloud modes, or mixed modes. The former two enjoy a single classically allowed region, with TS modes localized within the ISCO while cloud modes are located outside of it. Mixed modes feel two classically allowed regions separated by a centrifugal barrier. In the limit $\mu r_{b}\gg 1$, the cloud modes are dictated by the behavior of stable circular geodesics, as discussed above. The addition of KK momentum enriches the mode structure, since it introduces additional potential barriers, as discussed next.
@@
normal modes of TSs can carry KK momentum, $k=p/R_{y}$, where $p$ is an integer and $R_{y}$ the radius of the extra dimension. Far from the star, this yields the usual tower of mass corrections to the mode. Closer to the star, though, KK momentum results in a potential barrier that prevents the mode from reaching the star’s center. In fact, for large enough momentum modes cannot remain bound to the star and are radiated to infinity. Precisely, we find that a KK mode can be bound only if (taking $p>0$ without loss of generality)
@@
where $\lambda=1/\mu$ is the field’s Compton wavelength. Modes with $p\geq p_{\star}$ are necessarily radiated to infinity, so $p_{\star}$ defines the KK ionization threshold. Assuming that (2) holds, we find that, similarly to the zero-KK case, in the regime $\mu r_{b}\ll 1$ the spectrum becomes hydrogenic,
@@
where $P$ is the TS’s monopole charge. While the classification of neutral fields, based on their radial distribution, extends similarly to charged modes, the angular distribution of the latter changes significantly. In particular, charged modes can become localized at the north or south axis of the star, forming an axially-symmetric configuration, see Figure 2. We show that these states carry more angular momentum than their neutral counterparts, when normalized to have the same energy. This yields a gravitating version of a Thomson dipole in the strong gravity regime. [^1] We recall that a Thomson dipole is the system formed by an electric and a magnetic charge Thomson (2009). It constitutes an alternative to previous constructions where one attempts to form a gravitating Thomson dipole with BHs and point particles Garfinkle and Rey (1991); Bunster and Henneaux (2007); Kim and Moon (2007); Dyson and Pereñiguez (2023), and, it might possess a nonlinear completion similar to that of the superradiant threshold modes of magnetic BHs Pereñiguez et al. (2024); Cunha et al. (2025).

### v2

@@ Abstract
We study the bound states of a massive scalar field around a topological star, and show that these are strictly normal modes. This yields a genuine gravitational atom, sharply distinguishing horizonless objects from black holes. We show that the modes are controlled by the field’s Compton wavelength compared to the size of the star. When the Compton wavelength is large, the field forms a cloud with a hydrogen-like spectrum, while in the opposite regime it is localized along timelike trajectories. When the two scales are comparable the spectrum becomes richer, and we characterize it in detail allowing the field to carry electric charge and Kaluza–Klein momentum.
@@ I Introduction
The existence of event horizons in our universe is a robust prediction of general relativity (GR), as follows from combining singularity theorems Penrose (1965), cosmic censorship conjectures Penrose (1969) and the genericity of trapped surfaces Schoen and Yau (1983); Christodoulou (2008). Gravitational wave (GW) observatories Abbott et al. (2016, 2021) and long baseline interferometry Akiyama et al. (2019); Abuter et al. (2020) give support to such predictions on the empirical side. However, horizons pose major conceptual challenges such as the formation of spacetime singularities, the entropy problem and the information paradox Bena et al. (2022); Carballo-Rubio et al. (2025), none of them being resolvable by GR on its own. These reasons are strong enough to motivate the exploration of alternatives to the black hole (BH) paradigm, despite the strong theoretical and observational evidence for their existence Cardoso and Pani (2019); Berti et al. (2025).
@@
Topological solitons, as we will refer to the aforementioned objects in this work, are dramatically different from BHs near their surface. The latter feature an event horizon, which for an infalling observer is nothing but pure vacuum, yet it marks the point of no return for any such probe – rendering the inner boundary of a BH spacetime perfectly absorbing. Topological solitons, on the other hand, replace the horizon with a smooth end of spacetime – a “cap”, which, furthermore, supports non-negligible horizon-scale microstructure. This results in an inner boundary that is very much the opposite of the BHs of GR: classically the cap is perfectly reflecting, and its physics are certainly not of an empty spacetime. Such a dramatic change should surely be expressed in certain properties that topological solitons have or lack in contrast to BHs. Our aim here is to elucidate one such feature.
However, given these stark differences, how is it that topological solitons provide a good fit to the criteria for BH mimickers exhibited earlier? Most of the GW observations up to date are a consequence of the physics at the unstable photon sphere in the exterior of the spacetime, which effectively governs the linear response to perturbations at early times, irrespective of the presence or absence of a horizon Cardoso et al. (2016); Cardoso and Pani (2017). The topological solitons we are interested in all possess such a photon sphere, and, in specific cases, it has indeed been shown that the prompt response as well as the quasinormal modes (QNMs) relative to neutral, massless (or massive but radiative) fields exhibit a spectrum that parallels the one of BHs, although other sectors with new properties also emerge Heidmann et al. (2023a); Bianchi et al. (2023); Dima et al. (2024, 2025b). In addition, particle motion as well as gravitational and electromagnetic fluctuations have been explored in Dima et al. (2024); Bena et al. (2024); Dima et al. (2025b); Melis et al. (2025); Bini and Di Russo (2025a); Bianchi et al. (2025b); Bini and Di Russo (2025b); Di Russo et al. (2025); Bianchi et al. (2025c), giving strong evidence for their linear stability and again demonstrating important similarities with BHs. Finally, with respect to their tidal deformability, topological solitons are strikingly similar to magnetic BHs, both exhibiting vanishing tidal Love numbers relative to neutral fields, but non-vanishing ones when charged probes are considered, as shown recently in Ref. Pereñiguez and Karnickis (2025).
@@
One such aspect that has remained largely unexplored concerns the bound states of massive fields around compact objects. In the case of BHs, it is well understood that such fields – for example, dark matter candidates – can grow spontaneously, forming clouds that are bound to the hole, a system known as the gravitational atom Damour et al. (1976); Zouros and Eardley (1979); Detweiler (1980a). The mechanism yielding these clouds is the superradiant amplification due to the hole’s rotation, so, by definition, such gravitational atoms are at most quasi-stationary states, and require that the central hole rotates (see Brito et al. (2015a) for a review). These systems have been widely studied in the literature and have been considered to test possible dark matter candidates Arvanitaki et al. (2010); Arvanitaki and Dubovsky (2011), to obtain photon mass bounds Pani et al. (2012); Baumann et al. (2019a), and have been proposed as new sources of gravitational Yoshino and Kodama (2014); Brito et al. (2015b); Arvanitaki et al. (2015) and electromagnetic radiation Rosa and Kephart (2018); Ikeda et al. (2019); Chen et al. (2020). Moreover, when studying binaries, the cloud’s presence can affect the orbital dynamics to varying extents, depending on the cloud’s self-gravity Ferreira et al. (2017); Bar et al. (2019); Amorim et al. (2019); Vicente and Cardoso (2022); Duque et al. (2024); Dyson et al. (2025); Tomaselli et al. (2024a, b, 2023), and tidal effects can further enrich the binary dynamics Baumann et al. (2019b); Berti et al. (2019); Cardoso et al. (2020); Zhang and Yang (2019); Baumann et al. (2020).
@@
In the case of topological solitons, the perfectly reflecting inner boundary condition, resulting from the smooth end of spacetime in their interior, would imply that the states of such massive fields are not only bound at infinity, but also experience no dissipation at the surface of the soliton. We expect then that topological solitons will display a normal mode spectrum in addition to their QNMs – a property that is unattainable for the BHs of GR. This clear difference between BHs and TSs in the response to massive perturbations undoubtedly opens an interesting new avenue for exploration in view of the vast amount of research that has followed from the discovery of superradiant clouds around BHs.
@@
In this work we confirm this expectation in the case of one of the simplest such geometries that is also one of the most widely studied ones: topological stars (TSs) Bah and Heidmann (2021a, c). These are a one-parameter family of static solutions to the Einstein–Maxwell theory in five spacetime dimensions. Far from the star, the geometry is that of a four-dimensional Minkowski space endowed with a compact, circular extra dimension, as in the traditional Kaluza–Klein set up. Closer to the star, the geometry is smooth and supported by electromagnetic fluxes. Despite their simplicity, TSs contain the fundamental ingredients of any microstate geometry: a horizonless, smooth end-of-space, supported by electromagnetic flux. We probe them with an electrically charged, massive scalar field minimally coupled to the Einstein–Maxwell theory. We provide a comprehensive characterization of its bound states, that is, the states that are exponentially suppressed at infinity. Among our results, we show that any such mode is strictly stationary, thus proving that TSs are linearly stable under such perturbations with a characteristic normal mode spectrum. In this sense, TSs yield a more genuine notion of a gravitational atom than BHs, since the latter are only quasi-stationary and, in fact, unstable. We show that, in certain regimes, TS gravitational atoms share similarities with their BH counterparts, but they are, in general, markedly different. In addition, the KK momentum and electric charge endow the space of states with a rich structure, mostly absent in the BH case, which we analyze in detail. At the end of the paper, we discuss possible formation mechanisms of TS gravitational atoms focusing on potential instabilities of different nature.
@@
While our work focuses on the simple TS, we believe that our results on the existence of normal modes should be widely applicable to any horizonless geometry with perfectly reflecting inner boundary conditions. Irrespective of whether generalizations are possible, we believe that simple topological solitons are excellent, UV-motivated toy models for exploring horizonless physics with the aim of learning valuable lessons for the future. Building a smooth, horizonless geometry that mimics Kerr BHs as closely as possible in our observations is currently a hard technical challenge, but no solid theoretical arguments to the contrary are known Chakraborty and Heidmann (2025). In light of this peak into the future, we would like to emphasize another property of topological solitons that is in definite contrast to BHs and requires more attention. As we clarified at the beginning of this article, classically, the end product of gravitational collapse for sufficiently dense energy distributions is always a BH. The dynamical formation of topological solitons is thus necessarily a process that requires physics beyond classical GR.
@@
This paper is structured as follows. First, in Section I.1 we summarize our results for the reader’s convenience. Next, in Section II we discuss generic aspects of massive bound states of TSs. This includes proving that bound modes are strictly normal and showing that in the long (short) Compton wavelength regime, modes are hydrogenic (governed by timelike geodesics). In Section III we provide a thorough analysis of TS normal modes, discussing their regime of existence, classifying them, and also computing their spectrum numerically, including the scenarios where the field carries KK momentum and is electrically charged. We conclude in Section IV with a discussion of our results and open questions.
I.1 Executive summary
@@
Our results fall into two classes: those that are expected to hold in a qualitatively similar way for compact objects other than TS, and those that are TS-specific. We summarize them here. We remind the reader that TSs are characterized by two length scales, the TS’s size $r_{b}$ and the length of the compact extra-dimension $R_{y}$.
@@ I.1.1 Generic results
considering neutral probes (no charge or KK momentum), in the limit $\mu r_{b}\gg 1$, we show that solutions are governed by congruences of timelike geodesics. This allows us to identify two broad families of modes, dubbed cloud and inner modes. The former have a large angular momentum $\ell\gg\mu r_{b}$ and are exponentially localized around stable circular geodesics. They cease to exist for angular momenta lower than $\ell\sim\sqrt{3}r_{s}\mu$, where $r_{s}$ is a length fixed by $r_{b}$ and $R_{y}$ and related to the ISCO by $r_{\rm ISCO}=3r_{s}$. The behavior of modes changes abruptly below this threshold and becomes dictated by the inner structure of the geometry: these are the inner modes. The separation between the two families, at $\ell\sim\sqrt{3}r_{s}\mu$, is associated with the ISCO, and the modes inherit some of its properties. This is illustrated in Fig. 1.
@@
in the opposite regime, where $\mu r_{b}\ll 1$, the normal mode spectrum (restricting attention to zero-KK modes) becomes hydrogenic,
$\omega_{n}\approx\mu\left(1-\frac{\mu^{2}r_{s}^{2}}{8}\frac{1}{n_{\text{H}}^{2}}\right)\,,\quad\left(n_{\text{H}}=1,2,...\,,\ \ n_{\text{H}}>\ell\right).$ (1)
A similar behavior is also present in the superradiant quasi-bound states of rotating BHs Detweiler (1980b); Baumann et al. (2019a), in the limit $\mu M\ll 1$, where $M$ is the BH’s mass. One difference is the fine structure constant, which here is proportional to $\mu r_{s}$, while it is proportional to $\sim\mu M$ for BHs. However, the most important difference is that in the TS case these modes are truly stationary states, unlike for BHs, where they are generically unstable Hod (2012); Herdeiro and Radu (2014); Carretero et al. (2025); Nicoules et al. (2025).
@@ I.1.2 TS-specific results
for probes with no KK momentum (yet possibly charged), the gravitational potential of TSs for bound modes can yield either one or two classically allowed regions where modes can be localized. This leads to a classification of bound states in TS modes, cloud modes, or mixed modes. The former two have a single classically allowed region, with TS modes localized within the ISCO while cloud modes are located outside of it. Mixed modes have two classically allowed regions separated by a centrifugal barrier. In the limit $\mu r_{b}\gg 1$, the cloud modes are dictated by the behavior of stable circular geodesics, as discussed above. The addition of KK momentum enriches the mode structure by introducing additional potential barriers, as discussed next.
@@
normal modes of TSs can carry KK momentum, $k=p/R_{y}$, where $p$ is an integer and $R_{y}$ the radius of the extra dimension. Far from the star, this yields the usual tower of mass corrections to the mode. Closer to the star, though, KK momentum results in a potential barrier that prevents the mode from reaching the star’s center. In fact, for large enough momentum modes cannot remain bound to the star and are radiated to infinity. More precisely, we find that a KK mode can be bound only if (taking $p>0$ without loss of generality)
@@
where $\lambda=1/\mu$ is the field’s Compton wavelength. Modes with $p\geq p_{\star}$ are necessarily radiated to infinity, so $p_{\star}$ defines the KK ionization threshold. Assuming that Eq. (2) holds, we find that, similarly to the zero-KK case, in the regime $\mu r_{b}\ll 1$ the spectrum becomes hydrogenic,
@@
where $P$ is the TS’s monopole charge. While the classification of neutral fields, based on their radial distribution, extends similarly to charged modes, the angular distribution of the latter changes significantly. In particular, charged modes can become localized at the north or south axis of the star, forming an axially symmetric configuration, see Fig. 2. We show that these states carry more angular momentum than their neutral counterparts, when normalized to have the same energy. This yields a gravitating version of a Thomson dipole in the strong gravity regime (recall that a Thomson dipole is the system formed by an electric and a magnetic charge Thomson (2009)). It constitutes an alternative to previous constructions where one attempts to form a gravitating Thomson dipole with BHs and point particles Garfinkle and Rey (1991); Bunster and Henneaux (2007); Kim and Moon (2007); Dyson and Pereñiguez (2023), and it might possess a nonlinear completion similar to that of the superradiant threshold modes of magnetic BHs Pereñiguez et al. (2024); Cunha et al. (2025).